OpenStudy (anonymous):

Line BD bisects angle ABC and angle ABC=4x+5, and angle ABD = 3x-1. What is the value of x?

OpenStudy (zale101):

First, do the equation like this 4x + 5 = 3x - 1 and solve for X

OpenStudy (mathstudent55):

 \( m<ABD = \dfrac{1}{2}m<ABC \) \(3x - 1 = \dfrac{1}{2}(4x + 5) \) \( 2(3x - 1) = 4x + 5\) Solve for x.